# Self-Guide: Better Task-Specific Instruction Following via Self-Synthetic Finetuning

### v1

@@ 1 Introduction
In our empirical evaluation of Self-Guide using on multiple tasks from NaturalInstructions V2 (Wang et al., 2022), applying Self-Guide to an already instruction-tuned model (Vicuna-1.5-7B, (Zheng et al., 2023)) yields an absolute performance improvement of 17.9 points of ROUGE-L for open-ended generation tasks and 14.6 points of accuracy on classification tasks, compared against the baseline of prompting the same model with the same prompt.
@@ 2.1 Data Generation
The prompt templates employed vary based on the task type—generation or classification. For generation tasks, a simple prompt template instructs the model to generate a new input following the provided examples. In contrast, in classification tasks, the model is directed to generate an input corresponding to a randomly chosen label from the available set, allowing for a more balanced label distribution in the generated examples. Further details on the prompt templates are provided in Section A.1.1.
@@
2.2 Quality Optimization Temperature and Rule-based Filters
@@
Length Filter: While the lengths of demonstrative examples may exhibit bias, we believe these examples already possess decent representativeness for a specific task in terms of length distribution. We assume that the lengths of demonstrative examples follow a normal distribution for a specific task, and since most data points of a normal distribution fall within two standard deviations from the mean, we stipulate that the lengths of generated examples’ inputs and outputs should also approximate a normal distribution with mean $\mu$ and variance $\sigma^{2}$ calculated from the demonstrative examples. Specifically, the lengths are required to be within the range $(\mu-2\sigma,\mu+2\sigma)$. Other attributes like semantic similarity of generated examples are computationally expensive and lack clear, intuitive definitions. Hence, we opt for the efficient and tractable length attribute under this normality assumption.
@@ 2.3 Quality Optimization: Parameter Tuning
We want Self-Guide to generate training data matching to the desired distribution specified by an instruction and examples. This requires optimizing various hyper-parameters on labeled data points, including the initial number of generated inputs, the temperature for input generation, the temperature for output generation, finetuning parameters like training epochs, and so on. To achieve this, we tune parameters on a held-out set of existing tasks (and corresponding instructions). We do so by searching parameters that maximize worst task performance, in order to identify parameters that are likely to be “good enough” for a broad set of tasks (Michel et al., 2021):
@@ 3 Experimental Setup
Avg Metric: Exact Match 33.1 45.3 47.7 14.6
@@ 3.2 Base Model
We selected Vicuna-7b-1.5 (Zheng et al., 2023; Wan et al., 2024) as the foundational model for input generation, output generation, and fine-tuning for two primary reasons: Firstly, having undergone fine-tuning on an extensive and diverse corpus of user conversations from ShareGPT, this model demonstrates robust cross-task instruction-following capabilities. However, despite its extensive training, its performance is not deemed satisfactory, as evidenced by its relatively low ranking of 73rd on the LMSYS Chatbot Arena Leaderboard (Chiang et al., 2024), prompting us to investigate opportunities for improvement.
We employ the same evaluation metrics as in the Super-NaturalInstructions benchmark, namely Exact Match for classification tasks and Rouge-L for generation tasks. Exact Match measures the exact string match between the predicted and ground truth labels, while Rouge-L computes the longest common subsequence between the generated and reference texts. For a fair comparison, we use identical settings for our base model and Self-Guided model, including model size, batch size, decoding strategy (greedy search), and prompt template in  Section A.1.2. During inference, we leverage the efficient VLLM framework proposed by Kwon et al. (2023). For supervised finetuning with teacher forcing, we employ the Hugging Face TRL library (von Werra et al., 2020) due to its flexibility and ease of use for transformer-based language models.
@@ 3.3 Baselines
    Few-Shot ICL: As our primary baseline, we compare against directly prompting the LM. This directly relies on the inherent instruction-following abilities of the model.
@@ 4 Results and Analysis
Our main experiment results are shown in Table 1. Specifically, we report an absolute improvement of 14.6% for classification tasks and 17.9% for generation tasks in the benchmark’s metrics. Self-Guide demonstrates its effectiveness in guiding LLMs toward task-specific expertise, even in extremely data-limited situations. This highlights the potential of self-synthesized data in adapting LLMs for specialized tasks from a more scalable perspective.
@@
Avg 52.8 59.4 41.5 45.2 59.3 27
Table 2: We compare in-context learning and finetuning on their ability to learn either from 3 manually written demonstrations or from a varying number of synthetically generated demonstrations. We find finetuning is better at leveraging synthetic data than in-context learning. Finetuning on self-synthesized data also gives better performance than finetuning on gold few-shot examples.
@@ 4.1 Comparing self-synthesized examples with gold few-shot examples
In  Table 2, we see that using generated training data with finetuning improves over 3-shot learning on average for both classification and generation tasks (comparing the “Self-Guide” and “3-Shot Finetuning” columns). On the other hand, comparing the “3-Shot ICL” and “Self-ICL” columns, we see that using generated training data with in-context learning helps only for generation tasks; for classification tasks, this strategy often leads to deteriorated performance. In Self-ICL for classification tasks, the prompted model tends to generate new input-output demonstrations rather than respond to the final input to be completed, leading to a significant performance decrease. We hypothesize that the model fails to segment the instruction from the demonstration examples, thereby failing to systematically classify the provided inputs according to the task definition.
@@ 4.2 Finetuning outperforms in-context learning on synthetic data
How much is the learning algorithm — finetuning versus in-context learning — responsible for the effectiveness of Self-Guide? To test this, we compare in-context learning on as many self-synthesized examples as can fit in the context window of an LM (“Self-ICL”) with finetuning on the same set of self-synthesized examples. We find in Table 2, compared to in-context learning, finetuning on the same generated examples achieves substantially better performance, with an average improvement of 14 absolute percentage points across almost all tasks. Interestingly, finetuning on just a few self-synthesized examples (e.g. 7) can yield considerable performance gains, which highlights the data efficiency of Self-Guide.
@@ 4.3 Self-Guide aligns LMs with the correct label distribution in many cases
To demonstrate the high quality of data generated by Self-Guide from a qualitative perspective, we analyze the distance between the output distributions of the models before and after Self-Guide and the ground truth distribution for classification tasks.
@@ 4.5 Noise filter is crucial for classification tasks, while length filter is crucial for generation task
Ablation studies results below in Table 5 found that removing the ablation filter decreases classification accuracy by 4.1% while removing the length filter decreases generation accuracy by 3.8%. We think that the ablation filter is crucial for classification tasks as it removes outputs with superfluous content, retaining only the labels. For generation tasks, the length filter enhances performance by excluding lengthy or too short responses, improving the Rouge-L score. These studies demonstrate the effectiveness of our rule-based filters.
@@ 5 Related Work
Manually crafted datasets have played a pivotal role in supervising and augmenting various NLP task systems (Rozière et al., 2023; Yuan et al., 2023). However, traditional manual annotation processes are often time-consuming, labor-intensive, costly, and non-scalable. Additionally, the possibility of human errors and the lack of domain expertise among annotators in complex tasks cannot be overlooked (Braylan & Lease, 2021; Kang et al., 2023). Recent works have tried to alleviate the human labor associated with training data collection by utilizing the power of stronger models as data generators (Patel et al., 2024; Song et al., 2024). These methods can be broadly categorized based on their reliance on seed data, the need for human curation, and stronger teacher LLMs. Viswanathan et al. (2023) proposed Prompt2Model, which combines retrieval methods and distillation to directly gather training data from external sources without seed data. Zhou et al. (2023) introduced LIMA, which generates possible responses using LLMs given existing instructions, and then curates them with human annotators. Li et al. (2023) automated the generation of instructions through back-translation on large external corpora, eliminating the need for seed data (Self-Alignment). However, these methods still require more powerful models or seed datasets to cultivate high-quality examples. Our work investigates the self-generation abilities of LLMs as a means of reducing the need for external resources compared to previous efforts.
@@ 7 Limitations and Ethical Considerations
The primary limitation of our work is that we focus all of our experiments and models on English NLP tasks. Language technologies already promote systematic inequalities between languages and communities (Blasi et al., 2021); Self-Guide brings the possibility of exacerbating these inequalities. On the other hand, Self-Guide also offers the tempting potential of improving LMs’ ability to execute instructions specified in non-English languages to empower non-English LLM-based Agents (Chen et al., 2024). We consider this to be an essential direction for future work, and we are actively pursuing this.
@@ 8 Acknowledgments
This work used computing resources provided by the Microsoft Accelerate Foundation Models Research Program.
@@ A.1.1 Input Generation Prompt Template
These are some addtional [input]. Their formats and contents may not be accurate. However, you may also refer to the content of them.
@@
Afters seeing example inputs, generate a new [input]. Before generating the new [input], ensure that you strictly adhere to the rules of the new [instruction] and follow the format of high-quality [input].
@@
These are some addtional [input]. Their formats and contents may not be accurate. However, you may also refer to the content of them.
@@
Afters seeing example inputs, generate a new [input] for which the expected [output] is conditional_label. Before generating the new [input], ensure that you strictly adhere to the rules of the new [instruction] and follow the format of high-quality [input].
@@ A.2 Prompt Sensitivity Experiment
Raw Model Self-Guided Expert Model
@@
Classif. task190 27.0 21.1 19.8 7.1 66.7 66.7 66.7 0.0
task1529 8.5 11.0 29.5 21.0 61.6 54.9 53.1 8.4
task1612 51.3 50.9 48.8 2.4 42.6 42.4 42.8 0.4
task329 29.1 28.7 27.6 1.4 46.3 45.6 45.9 0.6
Avg 28.9 27.9 31.4 8.0 54.3 52.4 52.1 2.4
Generat. task281 46.8 43.4 32.4 14.3 52.3 46.9 47.3 5.4
task1195 49.9 44.1 57.3 13.1 83.7 85.4 85.5 1.7

### v2

@@ 1 Introduction
In our empirical evaluation of Self-Guide using on multiple tasks from Super-NaturalInstructions V2 (Wang et al., 2022), applying Self-Guide to an already instruction-tuned model (Vicuna-1.5-7B, (Zheng et al., 2023)) yields an absolute performance improvement of 17.9 points of ROUGE-L for open-ended generation tasks and 14.5 points of accuracy on classification tasks, compared against the baseline of prompting the same model with the same prompt.
@@ 2.1 Data Generation
The prompt templates employed vary based on the task type-generation or classification. For generation tasks, a simple prompt template instructs the model to generate a new input following the provided examples. In contrast, in classification tasks, the model is directed to generate an input corresponding to a randomly chosen label from the available set, allowing for a more balanced label distribution in the generated examples. Further details on the prompt templates are provided in Section A.1.1.
@@
2.2 Quality Optimization: Temperature and Rule-based Filters
@@
Length Filter: While the lengths of demonstrative examples may exhibit bias, we assume these examples already possess decent representativeness for a specific task in terms of length distribution. Based on this assumption, we further assume that the lengths of demonstrative examples follow a normal distribution for a specific task, and since most data points of a normal distribution fall within two standard deviations from the mean, we stipulate that the lengths of generated examples’ inputs and outputs should also approximate a normal distribution with mean $\mu$ and variance $\sigma^{2}$ calculated from the demonstrative examples. Specifically, the lengths are required to be within the range $(\mu-2\sigma,\mu+2\sigma)$. Other attributes like semantic similarity of generated examples are computationally expensive and lack clear, intuitive definitions. Hence, we opt for the efficient and tractable length attribute under this normality assumption.
@@ 2.3 Quality Optimization: Parameter Tuning
We want Self-Guide to generate training data matching to the desired distribution specified by an instruction and examples. This requires optimizing various hyper-parameters on labeled data points, including the initial number of generated inputs, the temperature for input generation, the temperature for output generation, finetuning parameters like training epochs, and so on. To achieve this, we tune the parameters on a set of existing tasks and their corresponding instructions, and then evaluate the model on a held-out dataset. We do so by searching parameters that maximize worst task performance, in order to identify parameters that are likely to be “good enough” for a broad set of tasks (Michel et al., 2021):
@@ 3 Experimental Setup
Avg Metric: Exact Match 33.2 45.0 47.7 14.5
@@ 3.2 Base Model
We selected Vicuna-7b-1.5 (Zheng et al., 2023; Wan et al., 2024; Fan & Tao, 2024) as the foundational model for input generation, output generation, and fine-tuning for two primary reasons: Firstly, having undergone fine-tuning on an extensive and diverse corpus of user conversations from ShareGPT, this model demonstrates robust cross-task instruction-following capabilities. However, despite its extensive training, its performance is not deemed satisfactory, as evidenced by its relatively low ranking of 73rd on the LMSYS Chatbot Arena Leaderboard (Chiang et al., 2024), prompting us to investigate opportunities for improvement.
We employ the same evaluation metrics as in the Super-NaturalInstructions V2 benchmark, namely Exact Match for classification tasks and ROUGE-L for generation tasks. Exact Match measures the exact string match between the predicted and ground truth labels, while ROUGE-L computes the longest common subsequence between the generated and reference texts. For a fair comparison, we use identical settings for our base model and Self-Guided model, including model size, batch size, decoding strategy (greedy search), and prompt template in  Section A.1.2. During inference, we leverage the efficient VLLM framework proposed by Kwon et al. (2023). For supervised finetuning with teacher forcing, we employ the Hugging Face TRL library (von Werra et al., 2020) due to its flexibility and ease of use for transformer-based language models.
@@ 3.3 Baselines
    Few-Shot ICL (Prompting): As our primary baseline, we compare against directly prompting the LM. This directly relies on the inherent instruction-following abilities of the model.
@@ 4 Results and Analysis
Our main experiment results are shown in Table 1. Specifically, we report an absolute improvement of 14.5% for classification tasks and 17.9% for generation tasks in the benchmark’s metrics. Self-Guide demonstrates its effectiveness in guiding LLMs toward task-specific expertise, even in extremely data-limited situations. This highlights the potential of self-synthesized data in adapting LLMs for specialized tasks from a more scalable perspective.
@@
Avg 52.9 59.4 41.6 45.2 59.4 27
Table 2: We compare in-context learning and finetuning on their ability to learn either from 3 manually written demonstrations or from a varying number of synthetically generated demonstrations. We find finetuning is better at leveraging synthetic data than in-context learning. When comparing Self-Guide with 3-Shot Finetuning, finetuning on self-synthesized data yields better performance than finetuning on gold-standard few-shot examples.
@@ 4.1 Comparing self-synthesized examples with gold few-shot examples
In  Table 2, we see that using generated training data with finetuning improves over 3-shot finetuning on average for both classification and generation tasks (comparing the “Self-Guide” and “3-Shot Finetuning” columns). On the other hand, comparing the “3-Shot ICL” and “Self-ICL” columns, we see that using generated training data with in-context learning helps only for generation tasks; for classification tasks, this strategy often leads to deteriorated performance. In Self-ICL for classification tasks, the prompted model tends to generate new input-output demonstrations rather than respond to the final input to be completed, leading to a significant performance decrease. We hypothesize that the model fails to segment the instruction from the demonstration examples, thereby failing to systematically classify the provided inputs according to the task definition.
@@ 4.2 Finetuning outperforms in-context learning on synthetic data
How much is the learning algorithm — finetuning versus in-context learning — responsible for the effectiveness of Self-Guide? To test this, we compare in-context learning on as many self-synthesized examples as can fit in the context window of an LM (“Self-ICL”) with finetuning on the same set of self-synthesized examples. We find in Table 2, compared to in-context learning, finetuning on the same generated examples achieves substantially better performance, with an average improvement of around 20 absolute percentage points across almost all tasks. Interestingly, finetuning on just a few self-synthesized examples (e.g. 7) can yield considerable performance gains, which highlights the data efficiency of Self-Guide.
@@ 4.3 Self-Guide aligns LMs with the correct label distribution in many cases
To demonstrate the high quality of data generated by Self-Guide from a quantitative perspective, we analyze the distance between the output distributions of the models before and after Self-Guide and the ground truth distribution for classification tasks.
@@ 4.5 Noise filter is crucial for classification tasks, while length filter is crucial for generation task
Ablation studies results below in Table 5 found that removing the ablation filter decreases classification accuracy by 4.1% while removing the length filter decreases generation accuracy by 3.7%. We think that the ablation filter is crucial for classification tasks as it removes outputs with superfluous content, retaining only the labels. For generation tasks, the length filter enhances performance by excluding lengthy or too short responses, improving the ROUGE-L score. These studies demonstrate the effectiveness of our rule-based filters.
@@ 5 Related Work
Manually crafted datasets have played a pivotal role in supervising and augmenting various NLP task systems (Rozière et al., 2023; Yuan et al., 2023). However, traditional manual annotation processes are often time-consuming, labor-intensive, costly, and non-scalable. Additionally, the possibility of human errors and the lack of domain expertise among annotators in complex tasks cannot be overlooked (Braylan & Lease, 2021; Kang et al., 2023). Recent works have tried to alleviate the human labor associated with training data collection by utilizing the power of stronger models as data generators (Patel et al., 2024; Wang et al., 2024; Song et al., 2024). These methods can be broadly categorized based on their reliance on seed data, the need for human curation, and stronger teacher LLMs. Viswanathan et al. (2023); Ge et al. (2024) combine retrieval methods and distillation to directly gather training data from external sources without seed data. Zhou et al. (2023) introduced LIMA, which generates possible responses using LLMs given existing instructions, and then curates them with human annotators. Li et al. (2023) automated the generation of instructions through back-translation on large external corpora, eliminating the need for seed data (Self-Alignment). However, these methods still require more powerful models or seed datasets to cultivate high-quality examples. Our work investigates the self-generation abilities of LLMs as a means of reducing the need for external resources compared to previous efforts.
@@ 7 Limitations and Ethical Considerations
The primary limitation of our work is that we focus all of our experiments and models on English NLP tasks. Language technologies already promote systematic inequalities between languages and communities (Blasi et al., 2021; Zhao & Zhang, 2024; Zhao et al., 2024); Self-Guide brings the possibility of exacerbating these inequalities. On the other hand, Self-Guide also offers the tempting potential of improving LMs’ ability to execute instructions specified in non-English languages to empower trustworthy non-English LLM-based Agents (Chen et al., 2024; Sun et al., 2024; Huang et al., 2023). We consider this to be an essential direction for future work, and we are actively pursuing this.
@@ 8 Acknowledgments
This work was supported by the computing resources provided by the Microsoft Accelerate Foundation Models Research Program and the Tencent AI Lab Rhino-Bird Gift Fund. The authors would like to thank the reviewers for their valuable feedback. Additionally, we extend our gratitude to Professor Pengfei Liu at Shanghai Jiaotong University for his assistance with this work.
@@ A.1.1 Input Generation Prompt Template
These are some additional [input]. Their formats and contents may not be accurate. However, you may also refer to the content of them.
@@
After seeing example inputs, generate a new [input]. Before generating the new [input], ensure that you strictly adhere to the rules of the new [instruction] and follow the format of high-quality [input].
@@
These are some additional [input]. Their formats and contents may not be accurate. However, you may also refer to the content of them.
@@
After seeing example inputs, generate a new [input] for which the expected [output] is conditional_label. Before generating the new [input], ensure that you strictly adhere to the rules of the new [instruction] and follow the format of high-quality [input].
@@ A.2 Prompt Sensitivity Experiment
Raw Model Self-Guide Expert Model
@@
Classif. task190 27.0 21.1 19.8 7.2 66.7 66.7 66.7 0.0
task1529 8.5 11.0 29.5 21.0 61.6 54.9 53.1 8.5
task1612 51.3 50.9 48.8 2.5 42.6 42.4 42.8 0.4
task329 29.1 28.7 27.6 1.5 46.3 45.6 45.9 0.7
Avg 29.0 27.9 31.4 8.1 54.3 52.4 52.1 2.4
Generat. task281 46.8 43.4 32.4 14.4 52.3 46.9 47.3 5.4
task1195 49.9 44.1 57.3 13.2 83.7 85.4 85.5 1.8